3logbase5(Y) - logbaseY(5) = 2 (solve for Y)
Answer by stanbon(75887)

3logbase5(Y) - logbaseY(5) = 2 (solve for Y)
--
3log5(y)-logy(5) = 2
---
log5(y^3)-logy(5) = 2
---
log(y^3)/log(5) - log(5)/log(y) = 2
---
multiply thru by log(5)(log(y)) to get:
log(y)*3log(y) -log(5)*log(5) = 2log(y)*(log(5))
3[log(y)]^2 -2log(5)*log(y) - [log(5)]^2 = 0
-----
This is a quadratic in log(y).
Let log(y) = "m"
3m^2 - log(25)m - (log(5))^2 = 0
---
Solve for "m":
---
Graphing I get m = -0.233
----
So, log(y) = -0.233
Then y = 10^(-0.233) = 0.5848
=================================
Cheers,
Stan H.
==========================